A first order reaction is run with a catalyst at room temperature. A single data point is taken, showing >99% conversion at 5 hours. The catalyst is then reused, showing 97% conversion at 5 hours. The authors' conclusion is that the catalyst has retained its activity with reuse'. By how much could have their catalyst rate actually changed and they would still have the observed conversions?
Answer #1

every 2 % of conversion takes total time = 10 hours

therefore, 1% = 10/2 = 5 hours

hence , 100% = 10/2 x 100 = 500 hours

so, 100 % of catalyst conversion takes 500 hours